# Latent Action Pretraining From Videos

### v1

@@ Abstract
We introduce Latent Action Pretraining for general Action models (LAPA), an unsupervised method for pretraining Vision-Language-Action (VLA) models without ground-truth robot action labels. Existing Vision-Language-Action models require action labels typically collected by human teleoperators during pretraining, which significantly limits possible data sources and scale. In this work, we propose a method to learn from internet-scale videos that do not have robot action labels. We first train an action quantization model leveraging VQ-VAE-based objective to learn discrete latent actions between image frames, then pretrain a latent VLA model to predict these latent actions from observations and task descriptions, and finally finetune the VLA on small-scale robot manipulation data to map from latent to robot actions. Experimental results demonstrate that our method significantly outperforms existing techniques that train robot manipulation policies from large-scale videos. Furthermore, it outperforms the state-of-the-art VLA model trained with robotic action labels on real-world manipulation tasks that require language conditioning, generalization to unseen objects, and semantic generalization to unseen instructions. Training only on human manipulation videos also shows positive transfer, opening up the potential for leveraging web-scale data for robotics foundation model. We open-source the model checkpoints and code at latentactionpretraining.github.io.
@@ 1 Introduction
Vision-Language-Action Models (VLA) for robotics (Brohan et al., 2023; Kim et al., 2024) are trained by aligning large language models with vision encoders, and then finetuning it on on diverse robot datasets (Collaboration et al., 2023); this enables generalization to novel instructions, unseen objects, and distribution shifts (Michał et al., 2024). However, diverse real-world robot datasets mostly require human teleoperation, which makes scaling difficult. Internet video data, on the other hand, offers abundant examples of human behavior and physical interactions at scale, presenting a promising approach to overcome the limitations of small, specialized robotic datasets (Yang et al., 2024c). However, it is challenging to learn from internet video data for two major challenges: first, much of the raw data on the web lacks explicit action labels; second, the data distribution from the web is fundamentally different from the embodiments and environments of typical robotic systems (McCarthy et al., 2024). We propose Latent Action Pretraining for General Action Models (LAPA), an unsupervised approach to pretraining a robotic foundation model without the need for ground-truth robot action labels (Figure 1).
LAPA has two pretraining stages, followed by a fine-tuning stage to map the latent actions to real robot actions. In the first pretraining stage, we use a VQ-VAE-based objective (Van Den Oord et al., 2017) to learn quantized latent actions between raw image frames. Analogous to Byte Pair Encoding (Sennrich et al., 2016) used for language modeling, this can be seen as learning to tokenize atomic actions without requiring predefined action priors (e.g., end-effector positions, joint positions). In the second stage, we perform behavior cloning by pretraining a Vision-Language Model to predict latent actions derived from the first stage based on video observations and task descriptions. Finally, we fine-tune the model on a small-scale robot manipulation dataset with robot actions to learn the mapping from the latent actions to robot actions. In this work, we refer to both the proposed method and the resulting VLA models as LAPA.
@@
We measure performance on diverse manipulation videos, including existing robot video datasets (without utilizing ground-truth actions) and human manipulation datasets. Our results show that the proposed method significantly outperforms baseline methods of training manipulation policies from actionless videos, particularly in cross-environment and cross-embodiment scenarios. Furthermore, on real-world manipulation tasks, our method leads to a new monolithic VLA model, outperforming OpenVLA (Kim et al., 2024), the current state-of-the-art model Vision Language Action (VLA) model trained on a diverse mixture of datasets with ground-truth actions. These results demonstrate the effectiveness of learning unified quantized latent action representations across diverse robotic datasets featuring different embodiments (shown in Section 5.2). We further demonstrate that LAPA remains effective even when pretrained on only human manipulation video, outperforming models pretrained on Bridgev2 (Walke et al., 2023), one of the largest open-sourced robotic datasets. We observe that LAPA effectively captures environment-centric actions, including object and camera movements, which could be beneficial for downstream tasks like navigation or dynamic, non-quasistatic tasks. We expect that our method opens up the potential for building foundation models for robotics by pretraining on much larger web-scale video data.
@@
  • We propose Latent Action Pretraining for general Action models (LAPA), an unsupervised approach to pretraining a robotic foundation model to encode robotic skills from web-scale video data.
  • Experiments on simulation and real-world robot tasks show that our method not only significantly outperforms baseline methods for training robotic manipulation policies from actionless video, but also leads to a VLA model that outperforms the current state-of-the-art VLA model trained with ground-truth actions (by +6.22%), while achieving over 30x greater pretraining efficiency.
  • With LAPA, we qualitatively demonstrate that it is possible to use the learned world and action prediction models to simulate full trajectories in diverse environments, effectively building a neural simulation capable of performing closed-loop evaluations entirely through neural inference.
@@ 2 Related Work
2.1 Vision-Language-Action Models
Vision-Language Models (VLMs), which are trained on extensive internet-scale datasets encompassing text, image, and video, have demonstrated the ability to understand and generate both text and multimodal data (Liu et al., 2023; Team, 2024; Liu et al., 2024; Abdin et al., 2024). Building on the generalization power of VLMs, recent work has introduced Vision-Language-Action Models (VLAs), which extend VLMs by fine-tuning them on robotic action data to improve physical grounding (Brohan et al., 2023; Kim et al., 2024; Team et al., 2024; Collaboration et al., 2023). Recently, incorporating auxiliary objectives, such as visual traces (Niu et al., 2024), language reasoning paths (Michał et al., 2024), or constructing a conversational-style instruction dataset using robot trajectory data (Li et al., 2024a) during VLA training has been shown to improve performance. However, these methods still heavily rely on labeled action data, which limits the scalability of developing general VLAs, since they will be bounded by the amount of robotic data made available through human teleoperation. In contrast, our approach only requires labeled actions for finetuning, significantly reducing the burden of requiring a vast amount of human teleoperated robot data.
@@ 2.1 Vision-Language-Action Models
Videos contain rich information about dynamics and behavior which can be potentially beneficial for robot learning. However, most raw videos do not contain any action labels (McCarthy et al., 2024). One line of work attempts to leverage egocentric human videos (Grauman et al., 2022) to improve visual representations for visuomotor robotic tasks (Nair et al., 2022; Dasari et al., 2023). Some recent works explore the benefits of video generative models pretrained on human videos for downstream robotic tasks (Wu et al., 2024; Liang et al., 2024). Another line of work aims to learn useful information from human videos by learning from interactions (Zeng et al., 2024), affordances (Bahl et al., 2023; Kannan et al., 2023; Srirama et al., 2024; Shaw et al., 2023), or visual traces (Wen et al., 2023; Bharadhwaj et al., 2024b) extracted from human videos. However, all these works learn useful visual priors whereas our approach allows learning the mapping directly from perception to control during pretraining.
Another line of work aims to learn robot manipulation policies by retargeting human motions to robot motions. These works rely on off-the-shelf models such as hand pose estimators (Wang et al., 2023; Zhu et al., 2024; Shaw et al., 2023; Bharadhwaj et al., 2023; Ye et al., 2023; Qin et al., 2022) or motion capture systems (Yang et al., 2024a) to retarget the human motions directly to robot motions. However, these works either learn only task-specific policies or require large in-domain perfectly aligned human-robot data and do not leverage the diverse, web-scale human motion data to train a single generalist policy.
The last line of work assumes one has access to a small-scaled robotic dataset with action labels to train an inverse dynamic model (IDM), optical flow, or reinforcement learning models that predicts actions from future state rollouts generated by world models (Du et al., 2023; Ko et al., 2024; Yang et al., 2024b; Bharadhwaj et al., 2024a) trained on diverse web-scale human videos. Another related work (Baker et al., 2022) trains an IDM with a small amount of action-labeled data and uses the IDM to label large-scaled actionless web data and pretrain a visuomotor robot policy. We consider these methods as direct baselines to our proposed methods since they are scalable methods with the same experimental assumptions and learn both visual and control priors to train generalist robot policies.
@@ 3 LAPA: Latent Action Pretraining for general Action models
Figure 2: Overview of LAPA. (1) Latent Action Quantization: We first learn discrete latent actions in a fully unsupervised manner using the VQ-VAE objective. (2) Latent Pretraining: The VLM is trained to predict latent actions, essentially performing behavior cloning. After pretraining, we finetune the LAPA model on a small set of action-labeled trajectories to map the latent space to the end effector delta action space.
LAPA is divided into two stages: Latent Action Quantization and Latent Pretraining. First, we use a VQ-VAE based objective to capture the discretized latent delta information between consecutive frames in a video. Next, a pretrained VLM is trained to predict the latent action designated by the encoder of the Latent Action Quantization model, given the current image and the language instruction. After Latent Pretraining, we finetune the VLA model on a small number of ground-truth action-labeled trajectories to map the latent space to the actual action space. The overall process is illustrated in Figure 2.
@@ 3.1 Latent Action Quantization
To learn latent actions in a fully unsupervised manner, we train a latent action quantization model following Bruce et al. (2024) with a few modifications. Our latent action quantization model is an encoder-decoder architecture where the encoder takes the current frame $x_{t}$ and the future frame $x_{t+H}$ of a video with a fixed window size $H$ and outputs the latent action $z_{t}$ [^1] Although Bruce et al. (2024) conditioned on multiple past observations, we exclude previous frames due to computational constraints. We leave prepending past observations as future work.. The decoder is trained to take the latent action $z_{t}$ and $x_{t}$ and reconstruct $x_{t+H}$. Unlike Bruce et al. (2024), we use cross attention to attend $z_{t}$ given $x_{t}$ instead of additive embedding, which empirically leads to capturing more semantically meaningful latent actions. Our quantization model is a variant of C-ViViT tokenizer (Villegas et al., 2023) where the encoder includes both spatial and temporal transformer while the decoder only contains spatial transformer since our model uses only two image frames as input. Further model details are provided in Appendix A.
Our latent action quantization training model is based on the VQ-VAE objective (van den Oord et al., 2017). The VQ-VAE objective enables the latent action $z_{t}$ to be discrete tokens (codebooks), making it easy for VLMs to predict $z_{t}$. The latent action is represented using $s$ sequences from $|C|$ codebook vocabulary space. To avoid gradient collapse often observed in VQ-VAE, we utilize NSVQ (Vali & Bäckström, 2022) which replaces the vector quantization error to a product of original error and a normalized noise vector. We also apply codebook replacement technique from NSVQ during early training steps to maximize codebook utilization.
We utilize the encoder of our latent action quantization model as an inverse dynamics model in the next stage of LAPA and the decoder for generating neural-based closed-loop rollouts. Unlike previous works (Bruce et al., 2024; Valevski et al., 2024), LAPA trains both a world model that generates rollouts from the latent actions and a policy model that produces these latent actions through Latent Pretraining.
@@ 3.2 Latent Pretraining
We use the encoder of the latent action quantization model as an inverse dynamics model to label all $x_{t}$, given $x_{t+1}$, with $z_{t}$. Then, we pretrain a VLM to predict the $z_{t}$ given the language instruction of a video clip and the current image $x_{t}$. Instead of using the existing language model head of the VLM, we attach a separate latent action head of vocab size $|C|$. By default, we freeze only the vision encoder and unfreeze the language model during training. Since latent pretraining does not rely on ground truth actions, it opens the possibility of using any type of raw video paired with language instructions. Also, in contrast to traditional action granularity used in robotics (e.g. end-effector positions, joint positions, joint torques, etc.), our approach does not require any priors about the action hierarchy/granularity and is learned in an end-to-end manner simply by being optimized to best capture the ‘delta’ of consecutive observations in a given video dataset. We broadly refer to models having gone through latent pretraining as LAPA.
@@ 4 Experiments
In this section, we demonstrate the effectiveness of LAPA as a general-purpose pretaining method. Specifically, we focus on answering the following questions:
  • Q1. How does LAPA perform when there are cross-task, cross-environment, and cross-embodiment gaps between pretaining and fine-tuning?
  • Q2. Can LAPA learn superior priors compared to using ground-truth actions during pretraining in a multi-embodiment setting?
  • Q3. Can we create a performant LAPA solely from raw human manipulation videos?
@@ 4.1 Benchmarks and Environments
Table 1: Pretraining and fine-tuning dataset for each environment. Cross-Env denotes cross-environment, Cross-Emb denotes cross-embodiment, and Multi-Emb denotes multi-embodiment. For fine-tuning, MT denotes multi-task training and MI denotes tasks with diverse multi-instructions. Category denotes the main capability we are trying to quantify. Illustration of each environment is shown in Figure 3.
Environment Category Pretraining Fine-tuning
Dataset # Trajs Dataset # Trajs
LangTable In-Domain Sim (All 5 tasks) 181k 5 Tasks (MT, MI) 1k
Cross-Task Sim (All 5 tasks) 181k 1 Task (MI) 7k
Cross-Env Real (All 5 tasks) 442k 5 tasks (MT, MI) 1k
SIMPLER In-Domain Bridgev2 60k 4 Tasks (MT) 100
Cross-Emb Something v2 220k 4 Tasks (MT) 100
Real-World Cross-Emb Bridgev2 60k 3 tasks (MI) 450
Multi-Emb Open-X 970k 3 tasks (MI) 450
Cross-Emb Open-X 970k 1 task (MI, Bi-manual) 150
Cross-Emb Something v2 220k 3 tasks (MI) 450
(a) Language Table
(b) Simpler
(c) Real
Figure 3: Experimental Setups. (a) shows an example from the 440k real-world trajectories (top) and the 181k simulation trajectories (bottom) from the Language Table Benchmark. (b) shows the 4 different evaluation tasks we use with the SIMPLER environment. (c) shows the four different tasks that we perform in the real-world.
We evaluate the effectiveness of LAPA on 9 different task categories in 2 different simulation environments and 4 different real-world robotic tasks. Table 1 shows an overview of the pretraining and fine-tuning dataset for each setup and Figure 3 visualizes the simulation benchmark and real-world setups. More details of each evaluation setup are provided in Appendix B.
Language Table (Lynch et al., 2023) is a simulation where a robot performs 2 DOF actions to push blocks (see Figure 3) (a)). It includes 5 subtask categories: BlocktoBlock, BlocktoAbsolute, BlocktoBlockRelative, BlocktoRelative, and Separate. Each trajectory from both sim and real is classified to one of 87k behaviors specified via natural language instruction (e.g. ‘slide the blue cube down a bit’). During evaluation, we evaluate models for both seen and unseen scenarios, where unseen includes new objects (color and shape) and unseen combinations of seen objects.
SIMPLER (Li et al., 2024b) is a set of simulated environments for evaluating generalist robot manipulation policies. We assess our models on 4 tasks (Figure 3 (b)) using the 7 DOF WidowX robot arm. Since SIMPLER lacks fine-tuning trajectories, we collect 100 multi-task trajectories using successful rollouts from a VLA model trained on BridgeV2 data (Walke et al., 2023), with held-out trajectories differing in object orientation and position from the evaluation setup.
Real-World Tabletop Manipulation experiments used a 7 DOF Franka Emika Panda robot arm in three environments and 14 DOF bi-manual robot in a single environment (shown in Figure 3 (c)). We utilize three pretraining data sources: Bridgev2 (Walke et al., 2023), Open-X (Collaboration et al., 2023), and Something Something v2 (Goyal et al., 2017). Following Kim et al. (2024), for single arm experiments, we finetune on three multi-instruction tasks: (1) ‘Pick $<$object$>$ into Sink’, (2) ‘Cover $<$object$>$ with Towel’, and (3) ‘Knock $<$object$>$ Over’. For bi-manual robot experiments, the task is to ‘Put the $<$object1$>$ on the container and the $<$object2$>$ on the plate’. For all tasks, each task involves 150 trajectories across 15 objects. We use a task-specific partial success criterion for evaluation, following Kim et al. (2024).

### v2

@@ Abstract
We introduce Latent Action Pretraining, the first unsupervised method for pretraining Vision-Language-Action (VLA) models without ground-truth robot action labels. Existing Vision-Language-Action models require action labels typically collected by human teleoperators during pretraining, which significantly limits possible data sources and scale. In this work, we propose a method to learn from internet-scale videos that do not have robot action labels. We first train an action quantization model leveraging VQ-VAE-based objective to learn discrete latent actions between image frames, then pretrain a latent VLA model to predict these latent actions from observations and task descriptions, and finally finetune the VLA on small-scale robot manipulation data to map from latent to robot actions. Experimental results demonstrate that our method significantly outperforms existing techniques that train robot manipulation policies from large-scale videos. Furthermore, it outperforms the state-of-the-art VLA model trained with robotic action labels on real-world manipulation tasks that require language conditioning, generalization to unseen objects, and semantic generalization to unseen instructions. Training only on human manipulation videos also shows positive transfer, opening up the potential for leveraging web-scale data for robotics foundation models. We will open-source the model checkpoints and code at latentactionpretraining.github.io.
@@ 1 Introduction
Vision-Language-Action Models (VLA) for robotics (Brohan et al., 2023; Kim et al., 2024) are trained by aligning large language models with vision encoders, and then finetuning it on on diverse robot datasets (Collaboration et al., 2023); this enables generalization to novel instructions, unseen objects, and distribution shifts (Michał et al., 2024). However, diverse real-world robot datasets mostly require human teleoperation, which makes scaling difficult. Internet video data, on the other hand, offers abundant examples of human behavior and physical interactions at scale, presenting a promising approach to overcome the limitations of small, specialized robotic datasets (Yang et al., 2024c). However, it is challenging to learn from internet video data for two major challenges: first, much of the raw data on the web lacks explicit action labels; second, the data distribution from the web is fundamentally different from the embodiments and environments of typical robotic systems (McCarthy et al., 2024). We propose Latent Action Pretraining, an unsupervised approach to pretraining a robotic foundation model without the need for ground-truth robot action labels (Figure 1).
Latent Action Pretraining consists of two models that are learned sequentially, followed by a fine-tuning stage to map the latent actions to real robot actions. In the first pretraining stage, we use a VQ-VAE-based objective (Van Den Oord et al., 2017) to learn quantized latent actions between raw image frames. Analogous to Byte Pair Encoding (Sennrich et al., 2016) used for language modeling, this can be seen as learning to tokenize atomic actions without requiring predefined action priors (e.g., end-effector positions, joint positions). In the second stage, we perform behavior cloning by pretraining a Vision-Language Model to predict latent actions derived from the first stage based on video observations and task descriptions. Finally, we fine-tune the model on a small-scale robot manipulation dataset with robot actions to learn the mapping from the latent actions to robot actions. In this work, we refer to the resulting VLA models as LAPA.
@@
We measure performance on diverse manipulation videos, including existing robot video datasets (without utilizing ground-truth actions) and human manipulation datasets. Our results show that the proposed method significantly outperforms baseline methods of training manipulation policies without ground-truth action labels, particularly in cross-environment and cross-embodiment scenarios. Furthermore, on real-world manipulation tasks, our method leads to a new monolithic VLA model, outperforming OpenVLA, the current state-of-the-art model Vision Language Action (VLA) model trained on a diverse mixture of datasets with ground-truth actions. These results demonstrate the effectiveness of learning unified quantized latent action representations across diverse robotic datasets featuring different embodiments (shown in Section 5.2). We further demonstrate that Latent Action Pretraining remains effective even when pretrained on only human manipulation video, outperforming models pretrained on Bridgev2 (Walke et al., 2023), one of the largest open-sourced robotic datasets. We observe that LAPA effectively captures environment-centric actions, including object and camera movements, which could be beneficial for downstream tasks like navigation or dynamic, non-quasistatic tasks. We expect that our method opens up the potential for building foundation models for robotics by pretraining on much larger web-scale video data.
@@
  • We propose Latent Action Pretraining, an unsupervised approach to pretraining a robotic foundation model to encode robotic skills from web-scale video data.
  • Experiments on simulation and real-world robot tasks show that our method not only significantly outperforms baseline methods for training robotic manipulation policies that are pretrained without using ground truth action labels, but also leads to a VLA model that outperforms the current state-of-the-art VLA model trained with ground-truth actions (by +6.22%), while achieving over 30x greater pretraining efficiency.
  • We qualitatively demonstrate that it is possible to use LAPA as the action prediction model and decoder of the latent action quantization model as the world model by predicting future frames conditioned on the current observation and the latent action predicted by LAPA, effectively building a neural simulation capable of performing closed-loop evaluations entirely through neural inference.
@@ 2 Related Work
Vision-Language-Action Models
Vision-Language Models (VLMs), trained on large-scale internet datasets of text, images, and videos, have shown strong capabilities in understanding and generating both text and multimodal data (Liu et al., 2023; Team, 2024; Liu et al., 2024; Abdin et al., 2024). Leveraging this, recent advancements have introduced monolithic Vision-Language-Action Models (VLAs), which extend VLMs by fine-tuning them with robotic action data for enhanced physical grounding (Brohan et al., 2023; Kim et al., 2024; Team et al., 2024; Collaboration et al., 2023). Incorporating auxiliary objectives, such as visual traces (Niu et al., 2024), language reasoning paths (Michał et al., 2024), or creating conversational-style instruction datasets from robot trajectories (Li et al., 2024a), have further improved VLA performance. However, these methods remain dependent on labeled action data, limiting scalability. In contrast, our approach reduces reliance on human-teleoperated data by requiring labeled actions only for fine-tuning.
@@
Videos offer rich data for robot learning, but most lack action labels (McCarthy et al., 2024). Related work pretrains a vision encoder on egocentric human videos (Grauman et al., 2022) to improve visual representations (Nair et al., 2022; Dasari et al., 2023), or video generative models to generate future robot trajectories (Wu et al., 2024; Liang et al., 2024; He et al., 2024). Methods also extract diverse features from human videos such as interactions (Zeng et al., 2024), affordances (Bahl et al., 2023; Kannan et al., 2023; Srirama et al., 2024; Shaw et al., 2023), or visual traces (Wen et al., 2023; Bharadhwaj et al., 2024b). Some perform retargeting of human motions to robot actions to create robotic policies that involve hand pose estimators (Wang et al., 2023; Zhu et al., 2024; Shaw et al., 2023; Bharadhwaj et al., 2023; Ye et al., 2023; Qin et al., 2022) or motion capture systems (Yang et al., 2024a); these policies are usually task-specific or need aligned data human to robot data in the same environment. Finally, some train inverse dynamics models (IDMs), optical flow, or reinforcement learning models that predict actions from future state rollouts generated by world models (Du et al., 2023; Ko et al., 2024; Yang et al., 2024b; Bharadhwaj et al., 2024a) or use the IDM for active learning (Baker et al., 2022).
@@ 3 LAPA: Latent Action Pretraining for general Action models
Figure 2: Overview of Latent Action Pretraining. (1) Latent Action Quantization: We first learn discrete latent actions in a fully unsupervised manner using the VQ-VAE objective (Detail in Figure 8). (2) Latent Pretraining: The VLM is trained to predict latent actions, essentially performing behavior cloning. After pretraining, we finetune LAPA on a small set of action-labeled trajectories to map the latent space to the end effector delta action space.
Latent Action Pretraining consists of two models that are learned sequentially: Latent Action Quantization and Latent Pretraining. The overall process is illustrated in Figure 2. Note that we use the same pretraining dataset for Latent Action Quantization and Latent Pretraining.
@@ 3.1 Latent Action Quantization
To learn latent actions in a fully unsupervised manner, we train a latent action quantization model following Bruce et al. (2024) with a few modifications. Our latent action quantization model is an encoder-decoder architecture where the encoder takes the current frame $x_{t}$ and the future frame $x_{t+H}$ of a video with a fixed window size $H$ and outputs the latent action $z_{t}$ [^1] Although Bruce et al. (2024) conditioned on multiple past observations, we exclude previous frames due to computational constraints. We leave prepending past observations as future work.. The decoder is trained to take the latent action $z_{t}$ and $x_{t}$ and reconstruct $x_{t+H}$. Unlike Bruce et al. (2024), we use cross attention to attend $z_{t}$ given $x_{t}$ instead of additive embedding, which empirically leads to capturing more semantically meaningful latent actions. Our quantization model is a variant of C-ViViT tokenizer (Villegas et al., 2023) where the encoder includes both spatial and temporal transformer while the decoder only contains spatial transformer since our model uses only two image frames as input.
Our latent action quantization training model is based on the VQ-VAE objective (van den Oord et al., 2017), where the nearest quantized representation from the continuous embedding is retrieved from an embedding space where each embedding corresponds to a codebook. The VQ-VAE objective enables the latent action $z_{t}$ to be discrete tokens (codebooks), making it easy for VLMs to predict $z_{t}$. The latent action is represented using $s$ sequences from $|C|$ codebook vocabulary space. The sequence length $s$ is designated by the kernel size, stride and padding value of a CNN network which is used right before the vector quantization process. To avoid gradient collapse often observed in VQ-VAE, we utilize NSVQ (Vali & Bäckström, 2022) which replaces the vector quantization error to a product of original error and a normalized noise vector. We also apply stop gradient to the patch embedding of the frame $x_{t}$ during decoding to avoid representation collapse. Codebook replacement technique from NSVQ is applied during early training steps to maximize codebook utilization. Further model and training details are provided in Appendix A.
We utilize the encoder of our latent action quantization model as an inverse dynamics model in latent pretraining and the decoder for generating neural-based closed-loop rollouts. Unlike previous works (Bruce et al., 2024; Valevski et al., 2024), our approach trains both a world model that generates rollouts from the latent actions and a policy model that produces these latent actions through Latent Pretraining.
@@ 3.2 Latent Pretraining
We use the encoder of the latent action quantization model as an inverse dynamics model to label all frames $x_{t}$, given frame $x_{t+1}$, with latent action $z_{t}$. Then, we do action pretraining by using a pretrained VLM to predict the $z_{t}$ given the language instruction of a video clip and the current image $x_{t}$. Instead of using the existing language model head of the VLM, we attach a separate latent action head (a single MLP layer) of vocab size $|C|$. By default, we freeze only the vision encoder and unfreeze the language model during training. Since latent pretraining does not rely on ground truth actions, it opens the possibility of using any type of raw video paired with language instructions. Also, in contrast to traditional action granularity used in robotics (e.g. end-effector positions, joint positions, joint torques, etc.), our approach does not require any priors about the action hierarchy/granularity and is learned in an end-to-end manner simply by being optimized to best capture the ‘delta’ of consecutive observations in a given video dataset. We broadly refer to models having gone through latent pretraining as LAPA.
@@ 4 Experiments
In this section, we demonstrate the effectiveness of Latent Action Pretraining as a general-purpose pretaining method. Specifically, we focus on answering the following questions: Q1. How does LAPA perform when there are cross-task, cross-environment, and cross-embodiment gaps between pretaining and fine-tuning? Q2. Can LAPA learn superior priors compared to using ground-truth actions during pretraining in a multi-embodiment setting? Q3. Can we create a performant LAPA solely from raw human manipulation videos?
@@ 4.1 Benchmarks and Environments
We evaluate the effectiveness of LAPA on 9 different task categories in 2 different simulation environments and 3 different real-world robotic tasks. Table 3 shows an overview of the pretraining and fine-tuning dataset for each setup and Figure 9 in Appendix 3 visualizes the simulation benchmark and real-world setups. More details of each evaluation setup are provided in Appendix 3.
Language Table (Lynch et al., 2023) is a simulation where a robot performs 2 DOF actions to push blocks (see Figure 9) (a)). It includes 5 subtask categories: BlocktoBlock, BlocktoAbsolute, BlocktoBlockRelative, BlocktoRelative, and Separate. During evaluation, we evaluate models for both seen and unseen scenarios, where unseen includes new objects (color and shape) and unseen combinations of seen objects.
SIMPLER (Li et al., 2024b) is a set of simulated environments for evaluating generalist robot manipulation policies. We assess our models on 4 tasks (Figure 9 (b)) using the 7 DOF WidowX robot arm. Since SIMPLER lacks fine-tuning trajectories, we collect 100 multi-task trajectories using successful rollouts from a VLA model trained on BridgeV2 data (Walke et al., 2023).
Real-World Tabletop Manipulation experiments used a 7 DOF Franka Emika Panda robot arm in three environments (shown in Figure 9 (c)). We utilize three pretraining data sources: Bridgev2 (Walke et al., 2023), Open-X (Collaboration et al., 2023), and Something Something v2 (Goyal et al., 2017). Following Kim et al. (2024), we finetune on three multi-instruction tasks: (1) ‘Pick $<$object$>$ into Sink’, (2) ‘Cover $<$object$>$ with Towel’, and (3) ‘Knock $<$object$>$ Over’. Each task involves 150 trajectories across 15 objects. We use a task-specific partial success criterion for evaluation, following Kim et al. (2024).